# Skyfall-GS: Synthesizing Immersive 3D Urban Scenes from Satellite Imagery

### v1

@@ 3.2 Synthesize via Curriculum-learning Based Iterative Datasets Update
Multiple diffusion samples. While diffusion models effectively refine individual 3DGS renders, independently applying them across viewpoints introduces inconsistencies. Furthermore, 3DGS is well known to suffer from overfitting on single views, as pointed out by CoR-GS (Zhang et al., 2024b), causing artifacts when rendering from novel viewpoints.
@@
Ideally, the optimal denoising diffusion process should produce a distribution where all views maintain synchronized 3D appearance. However, independent 2D denoising on each view does not preserve 3D consistency, resulting in a denoising trajectory distribution that is a super-set of the optimal trajectories. Selecting a single denoising trajectory from this expanded distribution has negligible probability of yielding the optimal 3D-consistent result, leading to the artifacts observed in Figure˜8(c).
@@
To mitigate this, we synthesize $N_{s}$ independently refined samples per view, effectively sampling multiple trajectories from the denoising distribution. During optimization, the photometric loss $\mathcal{L}_{\text{color}}$ implicitly averages over these $N_{s}$ samples. Rather than committing to a single potentially suboptimal denoising path, this approach allows the 3DGS optimization to find a consensus representation that balances fidelity to individual samples while promoting geometric coherence across views. Ablation studies (Section˜4.2) and Figure˜8(c) confirm that this strategy successfully balances detail preservation with structural coherence.
@@
Iterative dataset update. Our curriculum-based Iterative Dataset Update (IDU), detailed in Algorithm 1, optimizes the 3DGS over $N_{e}$ episodes. In each episode, we render curriculum-guided views and refine them using FlowEdit (Kulikov et al., 2024) with specified prompts and strengths to generate a new training set. As the curriculum descends to lower altitudes, rendering quality steadily improves, particularly in previously occluded regions. We provide detailed parameters in Section˜A.1.
@@
6:$G$: Initial 3DGS from satellite-view training
@@
9:for i = 1 to $N_{e}$ do
10:  $\texttt{radius}\leftarrow R_{i}$
11:  $\texttt{elevation}\leftarrow E_{i}$
12:  $\texttt{cam\_views}\leftarrow\textsc{OrbitViews}(\{P\},\texttt{radius},\texttt{elevation},N_{v})$ $\triangleright$ Generate $N_{p}\times N_{v}$ views
13:  $\texttt{render\_views}\leftarrow\textsc{Render}(G^{\prime},\texttt{cam\_views})$ $\triangleright$ Render RGB images
14:  $\texttt{refine\_views}\leftarrow\textsc{FlowEditRefine}($render_views, $T_{\text{src}},T_{\text{tgt}},n_{\min},n_{\max},N_{s})$ $\triangleright$ Refine renders using FlowEdit
15:  $G^{\prime}\leftarrow\textsc{Train}(G^{\prime},\texttt{refine\_views})$ $\triangleright$ Update 3DGS using refined views
16:end for
17:return $G^{\prime}$
Optimization. For each episode $i$, we optimize the 3DGS using:
$\mathcal{L}_{\text{IDU}}(G_{i-1},\tilde{C}_{i})=\mathcal{L}_{\text{color}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\,\,$, (5)
where $G_{i-1}$ denotes the previous episode’s 3DGS model, and $\tilde{C}_{i}$ are the current refined images. We provide more implementation details in Section˜A.1.
4 Experiments
Datasets. We evaluate on high-resolution RGB satellite imagery from two sources. First, the 2019 IEEE GRSS Data Fusion Contest (DFC2019) (Le Saux et al., 2019) featuring WorldView-3 captures of Jacksonville, Florida ($\times2048$ pixels, 35 cm/pixel resolution). Camera parameters and sparse points were generated using SatelliteSfM (Zhang et al., 2019). We evaluate on four standard AOIs: JAX_004, JAX_068, JAX_214, and JAX_260, following Sat-NeRF (Marí et al., 2022) and EOGS (Savant Aira et al., 2025) protocols. Second, for geographic diversity, we use the GoogleEarth dataset (Xie et al., 2024) (training data for CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b)) containing NYC scenes. We use four scenes (004, 010, 219, 336) with training views rendered at an 80° elevation to approximate satellite conditions. Google Earth Studio (GES) (Google, 2024) renders serve as ground truth for both datasets. See Section˜A.2 for more detail about datasets.
Baselines. Our method connects satellite-based 3D reconstruction and city generation, requiring baselines from both fields. For satellite reconstruction, we compare with Sat-NeRF (Marí et al., 2022) and EOGS (Savant Aira et al., 2025) on DFC2019 (they require RPC input unavailable in GoogleEarth), plus Mip-Splatting (Yu et al., 2024) (enhanced with our appearance modeling) and CoR-GS Zhang et al. (2024b) on both datasets. [^1] Many methods lack available code or models: Sat2Scene (Li et al., 2024d), Sat2Vid (Li et al., 2021), EO-NeRF (Marí et al., 2023), Sat-DN (Liu et al., 2025b), SatelliteRF (Zhou et al., 2024b), Sat-Mesh (Qu & Deng, 2023), CrossViewDiff (Li et al., 2024c), SkySplat (Huang et al., 2025), and others. For city generation, we compare with CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b) on GoogleEarth (their training dataset). We use official implementations with default settings. All experiments run on a single RTX A6000 GPU.
Table 1: Quantitative comparison of different methods on DFC2019 (Le Saux et al., 2019). The results show that our method consistently achieves the best performance, indicating superior perceptual fidelity compared to all baselines. Metrics are computed between renders from each method and reference frames from GES.
Distribution Metrics Pixel-level Metrics*
@@ 4 Experiments
3D Reconstruction
Sat-NeRF (Marí et al., 2022) 88.36 4.868 10.05 0.269 0.864
EOGS (Savant Aira et al., 2025) 87.74 5.286 7.26 0.168 0.959
Mip-Splatting (Yu et al., 2024) 87.19 5.405 11.89 0.318 0.819
CoR-GS (Zhang et al., 2024b) 89.03 5.241 11.55 0.350 0.948
@@
Ours 27.35 2.086 12.38 0.321 0.791
Table 2: Quantitative comparison of different methods on GoogleEarth dataset (Xie et al., 2024). The results show that our approach consistently achieves the best performance, indicating superior perceptual fidelity compared to all baselines. Metrics are computed between renders from each method and reference frames from GES.
@@
Ours 9.91 2.009 14.28 0.298 0.394
Figure 6: Qualitative comparison on (a) DFC2019 and (b) GoogleEarth datasets. The leftmost column shows one representative example of the input satellite images. Our method outperforms all baselines in geometric accuracy and texture quality in low-altitude novel views, demonstrating enhanced building geometry, detailed facades, and reduced floating artifacts. Notably, our approach correctly preserves distinctive features such as the red pavement in scene 010 that competing methods miss. Unlike CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b), our method operates directly on satellite imagery without requiring pixel-aligned semantic maps or height-fields, enabling synthesis of complex geometric structures that more closely match GES references.
Evaluation metrics. We primarily use distribution-based metrics to quantify quality and diversity. We report $\text{FID}_{\text{CLIP}}$ (Kynkäänniemi et al., 2023) and CMMD (Jayasumana et al., 2024) that use the CLIP (Radford et al., 2021) backbone. This is based on their observations that the InceptionV3 (Szegedy et al., 2016) used in the classic FID (Heusel et al., 2017) and KID (Binkowski et al., 2018) is unsuitable for modern generative models. We complement these with user studies for perceptual quality assessment. We also report pixel-aligned metrics (PSNR (Huynh-Thu & Ghanbari, 2008), SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018)) as secondary references. While generally unsuitable for generative tasks, these metrics are meaningful for the Google Earth dataset, where all images come from the same consistent GES 3D representation, eliminating temporal variations.
4.1 Comparisons with Baselines
Quantitative comparison. We evaluate against both satellite reconstruction and city generation methods using distribution-based metrics. Evaluation images are created by dividing rendered frames into 144 patches ($512\times 512$ pixels). For comparison in the DFC2019 dataset, we render GES reference videos at 17° elevation, extracting 30 frames per AOI (4,320 images total). For comparison in the GoogleEarth dataset, we use $45^{\circ}$ elevation with 24 frames per scene (3,456 images total). We generate matching videos from all methods using identical camera parameters. Our method consistently outperforms all baselines across all metrics on both the DFC2019 and Google Earth datasets (Tables˜2 and 2), demonstrating effective reconstruction across diverse urban environments.
@@ 4.1 Comparisons with Baselines
Figure 7: User study results. Our method consistently outperforms Sat-NeRF (Marí et al., 2022), EOGS (Savant Aira et al., 2025), CoR-GS (Zhang et al., 2024b), CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b), achieving particularly high scores in geometric accuracy and overall perceptual quality. (a) details the comparison on the DFC2019 dataset (Le Saux et al., 2019), while subfigure (b) details the comparison on the GoogleEarth dataset (Xie et al., 2024).
Qualitative comparison. Figure 6(a) presents comparisons on the DFC2019 dataset against Sat-NeRF (Marí et al., 2022), EOGS (Savant Aira et al., 2025), and Mip-Splatting (Yu et al., 2024). All baselines exhibit significant distortions and blurry textures at lower viewpoints, while our baseline without IDU improves geometry but still shows floating artifacts and lacks facade detail. Our full approach achieves superior image quality. Figure 6(b) compares our approach on the GoogleEarth dataset against CityDreamer (Xie et al., 2024), GaussianCity (Xie et al., 2025b), and CoR-GS (Zhang et al., 2024b). While CityDreamer and GaussianCity generate plausible scenes, they produce oversimplified geometry and inaccurate textures, missing distinctive features such as the red pavement in scene 010 that our method correctly synthesizes. In contrast, our complete method achieves sharper building contours, enhanced texture fidelity, and reduced artifacts across both comparison scenarios. Notably, our approach successfully synthesizes plausible details for building facades occluded in the input satellite imagery and accurately reconstructs complex features including vegetation and multi-level architectures with finer surface details that better match the reference images. The visual quality approaches GES reference renders despite using only satellite imagery without ground-level data. Additional qualitative results are presented in Section˜A.2.
User studies. We conducted two comparative evaluations with 89 participants each: first, participants assessed the satellite input, GES reference video, Sat-NeRF, EOGS, CoR-GS, and our approach; second, participants compared the satellite input, GES reference video, CityDreamer, GaussianCity, CoR-GS, and our approach. Both studies evaluated geometric accuracy, spatial alignment, and overall quality, with full survey details in Section˜A.2. On the DFC2019 dataset, our method achieved dominant winrates of $\approx$97%/97%/97% vs. Sat-NeRF’s $\approx$3%/3%/3%, while EOGS and CoR-GS achieved 0%/0%/0%. On the GoogleEarth dataset, our approach maintained a clear advantage with $\approx$90%/90%/92% winrates vs. CityDreamer’s $\approx$4%/3%/3%, GaussianCity’s $\approx$3%/3%/3%, and CoR-GS’s $\approx$3%/4%/2%. These results consistently validate that our approach significantly outperforms all baselines under human perception across geometric accuracy, spatial alignment, and overall quality.
Rendering efficiency. Our method achieves 11 FPS on the modest NVIDIA T4 GPU, significantly outperforming CityDreamer’s 0.18 FPS despite running on the far more powerful NVIDIA A100, which offers 5× the CUDA cores and 10× the memory bandwidth. GaussianCity reaches comparable speeds (10.72 FPS) but requires the high-end A100. Furthermore, our fused representation enables real-time rendering at 40 FPS on consumer hardware (MacBook Air M2), demonstrating that our method enables high-quality 3D urban navigation without specialized computing resources.

### v2

@@ 3.2 Synthesize via Curriculum-Based Iterative Datasets Update
While diffusion models effectively refine individual 3DGS renders, independently applying the diffusion model across viewpoints introduces inconsistencies. Furthermore, 3DGS is well known to suffer from overfitting on single views, as pointed out by CoR-GS [114], causing artifacts when rendering from novel viewpoints.
@@
Ideally, the optimal denoising diffusion process should produce a distribution where all views maintain consistent 3D appearance. However, independent 2D denoising on each view does not preserve 3D consistency, resulting in a denoising trajectory distribution that is a superset of the optimal trajectories. Selecting a single denoising trajectory from this expanded distribution is unlikely to yield the optimal 3D-consistent result, leading to the artifacts observed in Figure˜9(c).
@@
To mitigate this, we synthesize $N_{s}$ independently refined samples per view, effectively sampling multiple trajectories from the denoising distribution. During optimization, the photometric loss $\mathcal{L}_{\text{color}}$ implicitly averages over these $N_{s}$ samples. Rather than committing to a single potentially suboptimal denoising path, this approach allows the 3DGS optimization to find a consensus representation that balances fidelity to individual samples while promoting geometric coherence across views. Ablation studies (Section˜4.2) and Figure˜9(c) confirm that this strategy successfully balances detail preservation with structural coherence.
3.2.4 Iterative dataset update.
@@
Our curriculum-based Iterative Dataset Update (IDU), detailed in Algorithm 1, optimizes the 3DGS over $N_{e}$ episodes. In each episode, we render curriculum-guided views and refine them using FlowEdit [36] with specified prompts and strengths to generate a new training set. As the curriculum descends to lower altitudes, rendering quality steadily improves, particularly in previously occluded regions, as illustrated in Figure˜6. We provide detailed parameters in the Supplementary.
@@
7:for i = 1 to $N_{e}$ do
8:  $\texttt{cam\_views}\leftarrow\textsc{OrbitViews}(\{P\},R_{i},E_{i},N_{v})$ $\triangleright$ Generate $N_{p}\times N_{v}$ views
9:  $\texttt{render\_views}\leftarrow\textsc{Render}(G^{\prime},\texttt{cam\_views})$ $\triangleright$ Render RGB images
10:  $\texttt{refine\_views}\leftarrow\textsc{Refine}(\texttt{render\_views},\Phi,N_{s})$ $\triangleright$ Refine renders using FlowEdit
11:  $G^{\prime}\leftarrow\textsc{Train}(G^{\prime},\texttt{refine\_views})$ $\triangleright$ Update 3DGS using refined views
12:end for
13:return $G^{\prime}$
Refer to caption Refer to caption Refer to caption
(a) After Reconstruction Stage (b) After Episode 1 (c) After Episode 2
Refer to caption Refer to caption Refer to caption
(d) After Episode 3 (e) After Episode 4 (f) After Episode 5 (Final)
Figure 6: Visualization of progressive refinement. This figure illustrates the step-by-step evolution of the synthesized 3D scene. Starting from the initial reconstruction state (a), the geometry and texture are progressively refined through successive stages of the iterative process (b-e), culminating in the final high-fidelity result (f).
3.2.5 Optimization.
For each episode $i$, we optimize the 3DGS using:
$\small\mathcal{L}_{\text{IDU}}(G_{i-1},\tilde{C}_{i})=\mathcal{L}_{\text{color}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\,\,$, (5)
where $G_{i-1}$ denotes the previous episode’s 3DGS model, and $\tilde{C}_{i}$ are the current refined images. We provide more implementation details in Supplementary.
4 Experiments
4.0.1 Implementation details.
The Reconstruction Stage runs for 30,000 iterations with $\lambda_{\text{D-SSIM}}=0.2$, $\lambda_{\text{op}}=10$, and $\lambda_{\text{depth}}=0.5$. The Synthesis Stage comprises $N_{e}=5$ episodes of 10,000 iterations each, with $N_{v}=6$ cameras and $N_{s}=2$ samples per look-at point. Training images are sampled 75% from IDU-refined views and 25% from original satellite views, preserving consistency with the input satellite imagery. All experiments run on a single RTX A6000 (48GB) GPU. Full details are in Supplementary.
4.0.2 Datasets.
We evaluate our method on high-resolution RGB satellite imagery from two sources. First, the 2019 IEEE GRSS Data Fusion Contest (DFC2019) [40] featuring WorldView-3 captures of Jacksonville, Florida ($\times2048$ pixels, 35 cm/pixel resolution). Camera parameters and sparse points are generated using SatelliteSfM [115]. We evaluate on four standard AOIs: JAX_004, JAX_068, JAX_214, and JAX_260, following Sat-NeRF [56] and EOGS [77]. Second, for geographic diversity, we use the GoogleEarth dataset [101] (training data for CityDreamer [101] and GaussianCity [103]) containing NYC scenes. We use four scenes (004, 010, 219, 336) with training views rendered at 80° elevation to approximate satellite conditions. Google Earth Studio (GES) [22] renders serve as ground truth for both datasets. See Supplementary for more detail.
4.0.3 Baselines.
As our method bridges satellite-based 3D reconstruction and city generation, we select baselines from both fields. For satellite reconstruction, we compare with Sat-NeRF [56] and EOGS [77] on DFC2019 (they require RPC input unavailable in GoogleEarth), plus Mip-Splatting [111] (with our appearance modeling enabled on DFC2019) and CoR-GS [114] on both datasets. For city generation, we compare with CityDreamer [101] and GaussianCity [103] on GoogleEarth dataset (their training dataset). [^1] Many methods lack available code or models: Sat2Scene [44] (inference only), Sat2Vid [45], EO-NeRF [57], Sat-DN [52], SatelliteRF [120], Sat-Mesh [72], CrossViewDiff [43], SkySplat [28], MagicCity [107], Sat3DGen [70], Sat2City [27], and Sat2RealCity [32]. We use official implementations with default settings.
4.0.4 Evaluation metrics.
Skyfall-GS primarily aims to synthesize and enhance the invisible or low-coverage regions from the satellite view; therefore, our evaluation focuses on quality and diversity assessment metrics. We report $\text{FID}_{\text{CLIP}}$ [37] and CMMD [30] which use the CLIP [73] backbone, as InceptionV3 [85] used in classic FID [25] and KID [4] is unsuitable for modern generative tasks. We complement these with user studies for perceptual quality assessment. We also report pixel-aligned metrics (PSNR [29], SSIM [94], LPIPS [117]) as secondary references. However, these metrics are unsuitable for generative tasks, as the synthetic elements in the obscured regions cannot match the invisible ground-truth. Moreover, on DFC2019, systematic illumination and color gaps between WorldView-3 imagery and GES references make pixel-level scores unreliable.
4.1 Comparisons with Baselines
Table 1: Quantitative comparison of different methods on DFC2019 [40]. The results show that our method consistently achieves the best performance, indicating superior perceptual fidelity compared to all baselines. Metrics are computed between renders from each method and reference frames from GES. Red, orange, and yellow indicate the best, second best, and third best results, respectively.
Distribution Metrics Pixel-level Metrics
[…]
3D Reconstruction
@@ 4.1 Comparisons with Baselines
[…]
@@
[…]
  • $\dagger$
    Enhanced with our appearance modeling.
Table 2: Quantitative comparison of different methods on GoogleEarth dataset [101]. The results show that our approach consistently achieves the best performance, indicating superior perceptual fidelity compared to all baselines. Metrics are computed between renders from each method and reference frames from GES. Red, orange, and yellow indicate the best, second best, and third best results, respectively.
@@
[…]
4.1.1 Quantitative comparison.
We evaluate against satellite reconstruction and city generation baselines by dividing rendered frames into 144 patches ($512\times 512$ pixels). Reference frames are extracted from GES at $17^{\circ}$ elevation for DFC2019 (30 frames/AOI, 4,320 total images) and $45^{\circ}$ for GoogleEarth (24 frames/scene, 3,456 total images); all methods are then evaluated on matching videos generated with identical camera parameters. As shown in Tables˜2 and 2, our method achieves highly competitive performance on both datasets. While Mip-Splatting attains a slightly better average LPIPS on GoogleEarth, our approach consistently outperforms all baselines across the vast majority of distribution and pixel-level metrics, a result further corroborated by the qualitative and user studies below, demonstrating robust synthesis across diverse urban environments. Notably, CoR-GS achieves a higher SSIM on DFC2019. We attribute this to its tendency to produce overly smooth and blurry reconstructions, which artificially inflate SSIM scores due to SSIM’s known insensitivity to blurring.
Figure 7: Qualitative comparison on (a) DFC2019 and (b) GoogleEarth datasets. The leftmost column shows a representative input satellite image. Our method consistently surpasses baselines in geometric accuracy and texture quality at low-altitude novel views. On (a), our synthesis exhibit sharp building contours and detailed rooftops, whereas Sat-NeRF [56] produces severe fog, EOGS [77] yields excessive darkness, CoR-GS [114] suffers from heavy blurring, and Mip-Splatting [111] exhibits prominent floaters and artifacts. On (b), CityDreamer [101] and GaussianCity [103] over-simplify building geometry and miss scene-specific details, CoR-GS [114] and Mip-Splatting [111] produce blurry facades, while our method delivers sharper geometry, richer facade details, and correctly recovers distinctive features such as the red pavement in scene 010.
4.1.2 Qualitative comparison.
Figure˜7(a) compares our method on DFC2019 against Sat-NeRF [56], EOGS [77], CoR-GS [114], and Mip-Splatting [111], all of which exhibit significant distortions and blurry textures at lower viewpoints. Figure˜7(b) further includes CityDreamer [101] and GaussianCity [103], which over-simplify geometry and miss scene-specific details (e.g., the red pavement in scene 010). Furthermore, our approach yields sharper building contours, higher texture fidelity, and fewer artifacts across both datasets. Notice that Skyfall-GS also recovers the challenging facade details in occluded regions and complex structures (e.g., vegetation and bridges) that require high precision. Additional qualitative results are provided in Supplementary.
@@
Figure 8: User study results. Our method consistently outperforms Sat-NeRF [56], EOGS [77], CoR-GS [114], Mip-Splatting [111], CityDreamer [101] and GaussianCity [103], achieving particularly high scores in geometric accuracy and overall perceptual quality. (a) details the comparison on the DFC2019 dataset [40], while (b) details the comparison on the GoogleEarth dataset [101].
4.1.3 User studies.
We conducted two user studies with 44 participants each, evaluating geometric accuracy, spatial alignment, and overall perceptual quality across 4 scenes per study. In each study, participants were presented with side-by-side renderings and asked to select the result that best matched each criterion. In the first study, participants compared Sat-NeRF [56], EOGS [77], CoR-GS [114], Mip-Splatting [111], and our approach; in the second, CityDreamer [101] and GaussianCity [103] replaced Sat-NeRF and EOGS (see Supplementary for full survey details). As shown in Figure˜8, our method achieves dominant win rates of $\approx$90–94% on DFC2019 and $\approx$79–82% on GoogleEarth across all three criteria, confirming strong and consistent human preference for our approach over all baselines in both satellite reconstruction and city generation settings.
Figure 9: Satellite-view training and IDU refinement ablation. (a) Opacity regularization reduces floating artifacts and yields denser reconstructions. (b) Pseudo-camera depth supervision improves geometry in planar, texture-less regions such as rooftops and roads. (c) Multiple diffusion samples per view reduce high-frequency geometric noise and enhance texture consistency, $N_{s}=2$ achieves the optimal results. (d) Curriculum learning progressively introduces challenging views, significantly improving geometric coherence in occluded regions over random sampling. (e) A context-free prompt causes only minor degradation in facade details, demonstrating robustness to prompts. Replacing our refinement with SDEdit leads to severe quality degradation, as noising-denoising fails to synthesize details while preserving satellite-defined geometry.
4.2 Ablation Studies
We conduct ablation studies on the JAX_068 AOI.
4.2.1 Ablation on the reconstruction stage.
We ablate appearance modeling, opacity regularization, and pseudo-camera depth supervision (see Table˜4 and Figure 9). For this ablation, we evaluate at higher elevation angles to assess render quality during the IDU process, rather than at the final low-altitude viewpoints. Appearance modeling is crucial for multi-date convergence, opacity regularization removes floating artifacts (Figure˜9(a)), and depth supervision sharpens geometry in planar regions (Figure˜9(b)). Together, they yield the lowest $\text{FID}_{\text{CLIP}}$/CMMD scores. Additionally, we validate geometric accuracy using DFC2019 LiDAR data [40] by unprojecting 3DGS depth renders into DSMs for comparison. Both opacity regularization and pseudo-depth supervision improve geometry, with their combination achieving the lowest MAE/RMSE.